OpenStudy (anonymous):

use the foil method (2x^5-6)(x^9-5)

OpenStudy (anonymous):

Do you know how to use the foil method?

Parth (parthkohli):

Wait

OpenStudy (anonymous):

Me or him?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay so a way to remember how to use the foil method is first outer, inner last.

Parth (parthkohli):

\(\begin{align} \text{F.O.I.L. = First Outside Inside Last}\\ \hline \\ &(ax+b)(cx+d)\\ \ \\ \text{First:}&(\color{blue}{ax}+b)(\color{blue}{cx}+d)\to \color{blue}{ax}\cdot\color{blue}{cx}=\color{blue}{acx^2}\\ \text{Outside:}&(\color{green}{ax}+b)(cx+\color{green}{d})\to \color{green}{ax}\cdot\color{green}{d}=\color{green}{adx}\\ \text{Inside:}&(ax+\color{red}{b})(\color{red}{cx}+d)\to \color{red}{b}\cdot\color{red}{cx}=\color{red}{bcx}\\ \text{Last:}&(ax+\color{orange}{b})(cx+\color{orange}{d})\to \color{orange}{b}\cdot\color{orange}{d}=\color{orange}{bd}\\ \hline \\ (ax+b)(cx+d)&=\color{blue}{acx^2}+\color{green}{adx}+\color{red}{bcx}+\color{orange}{bd} \end{align}\)

Parth (parthkohli):

Get it?

OpenStudy (anonymous):

i hate math lol

OpenStudy (compassionate):

\[2x^14 - 6x^9 - 10x^5 + 30\]

OpenStudy (compassionate):

That is two-x to the fourteenth power --- Minus six-x to the ninth, minus ten-x to the fifth, plus thirty.

OpenStudy (anonymous):

thanks